# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v1

@@ C.2 Differential Privacy
$\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}((B+\sigma\max\{1,\frac{\|G\|}{S}\})\sqrt{\frac{d}{m}})$
for all $i=1,\dots,B$ with high probability.
Appendix D Additional experimental details
@@
The architecture of our deep image prior network is structured on the foundation of PixelCNN++ as per (Salimans et al., 2017), and it utilizes a U-Net as depicted by (Ronneberger et al., 2015), which is built upon a Wide-ResNet according to (Zagoruyko and Komodakis, 2016), To simplify the implementation, we substituted weight normalization (Salimans and Kingma, 2016) with group normalization (Wu and He, 2018). The models operating on $32\times 32$ utilize four distinct feature map resolutions, ranging from $32\times 32$ to $4\times 4$. Each resolution level of the U-Net comprises two convolutional residual blocks along with self-attention blocks situated at the $16\times 16$ resolution between the convolutional blocks. When available, the reconstructed features are incorporated into each residual block.
@@
We adopt ResNet-18 (He et al., 2016) architecture as our backbone network. The network is pre-trained for one epoch for better results of gradient inversion. Note that this architecture results in a more challenging setting given its minimal information retention capacity and is observed to cause a slight drop in quantitative performance and large variance (Jeon et al., 2021). We further add an extra linear layer before the classification head for feature reconstruction. In order to satisfy the concentration bound $\tilde{O}(B\sqrt{\frac{d}{m}})$ of feature reconstruction, this linear layer has $512\times 32=16384$ neurons.
@@
The attack minimize the objective function given in Eq. 3. We search $\alpha_{\rm TV}$ in $\{0,0.001,0.005,0.01,0.05,0.1,0.5\}$ and $\alpha_{\rm BN}$ in $\{0,0.0005,0.001,0.01,0.05,0.01\}$ for attack without defense, and apply the best choices $\alpha_{\rm TV}=0.01$ and $\alpha_{\rm BN}=0.001$ for all defenses (we did not tune $\alpha_{\rm TV}$ and $\alpha_{\rm BN}$ for each defense as we observe the proposed attack method is robust in terms of these two hyper-parameters). We optimize the attack for 10,000 iterations using Adam (Kingma and Ba, 2014), with the initial learning rate set to $0.00001$. We decay the learning rate by a factor of $0.1$ at $3/8$, $5/8$, and $7/8$ of the optimization. For feature matching loss strength $\alpha_{f}$, we search in $\{0.01,0.05,0.1,0.5\}$ for each defense and find that $0.1$ is the optimal choice.
@@
Zhu et al. (2019); Geiping et al. (2020) suggests that small batch size is vital for the success of the attack. We therefore intentionally evaluate the attack with three small batch sizes $2$, $4$, and $8$ to test the upper bound of privacy leakage, and the minimum (and unrealistic) batch size $1$, and two small but realistic batch sizes $16$ and $32$.
@@
The feature reconstruction method suggested by Wang et al. (2023) always produces features that are normalized and may have a sign opposite to the actual features. To compare the similarity between the original and the reconstructed features, we use cosine similarity, as it doesn’t change with feature magnitude. We square this similarity to avoid any issues with mismatched signs.
When dealing with multiple features at once (batch size greater than 1), it’s not always obvious in which order they were reconstructed. This can tempt one to try out all possible orders to find the correct one, especially when errors might be intentionally induced. But we’ve used a simpler (but greedy) way. We compare each reconstructed feature with the original ones in sequence to figure out the correct order, ensuring each feature is only paired once. This prevents a situation where one reconstructed feature is incorrectly matched with several original features. Experimental results suggest this way works well.
Appendix E Additional experimental results

### v2

@@ E.1 Dropout
$R_{full}\geq\sqrt{\frac{1}{B}\frac{\sigma^{2}d^{2}B^{2}}{(1-p)\mathrm{tr}(J({\bm{x}})^{\top}J({\bm{x}}))}}=\tilde{\Omega}\left(\sigma\sqrt{\frac{d}{(1-p)m}}\right)$
where $J({\bm{x}})$ is the complete Jacobian matrix without deleted rows.
E.2 Gradient Clipping
Clipping reduces the size of the gradient when the gradient size is too large. The noisy gradient could be written as
$\min\left(\left\|\nabla_{\theta}\sum_{i=1}^{B}(y-f({\bm{x}}_{i}))^{2}\right\|,C\right)\frac{\nabla_{\theta}\sum_{i=1}^{B}(y-f({\bm{x}}_{i}))^{2}}{\left\|\nabla_{\theta}\sum_{i=1}^{B}(y-f({\bm{x}}_{i}))^{2}\right\|}+{\bm{\epsilon}},$
where $\theta$ stands for the model parameters and ${\bm{\epsilon}}$ is random noise with distribution $N(0,\frac{\sigma^{2}}{m^{2}}{\bm{I}}_{md+m})$.
When the norm of the gradient is smaller than the constant $C$, the lower bound remains unchanged. When the norm is higher than the constant $C$, the lower bound is of scale $\Omega(\frac{d\left\|\nabla_{\theta}\sum_{i=1}^{B}(y-f({\bm{x}}_{i}))^{2}\right\|^{2}}{mC^{2}})\sigma^{2}$. (When the number of model parameters is high we could ignore the effect of changes in the norm on the bound.)
Lemma E.1.
Under Assumption D.1, and for some constant $C$, when the norm of the model gradients $\left\|G\right\|\geq C$, with high probability the lower bound of minimax risk $R_{L}$ with the observation
$O(S)=C\frac{G}{\left\|G\right\|}+{\bm{\epsilon}}=C\frac{\nabla_{\theta}\sum_{i=1}^{B}(y-f({\bm{x}}_{i}))^{2}}{\left\|\nabla_{\theta}\sum_{i=1}^{B}(y-f({\bm{x}}_{i}))^{2}\right\|}+{\bm{\epsilon}}$
is of the scale $R_{L}\geq\tilde{\Omega}(\sigma\sqrt{\frac{d}{m}}\frac{\left\|G\right\|}{C})$, where ${\bm{\epsilon}}$ follows $\mathcal{N}(0,\sigma^{2}{\bm{I}}_{md+m})$ and $G=\nabla_{\theta}\sum_{i=1}^{B}(y-f({\bm{x}}_{i}))^{2}$.
Proof.
Denote ${\bm{g}}_{\theta}=C\frac{G}{\left\|G\right\|}$, then for each ${\bm{x}}_{k}$,
$\nabla_{{\bm{x}}_{k}}{\bm{g}}_{\theta}=$ $C\frac{\nabla_{{\bm{x}}_{k}}G}{\left\|G\right\|}+C\left(\nabla_{{\bm{x}}_{k}}\frac{1}{\left\|G\right\|}\right)G^{\top}$
$=$ $C\frac{\nabla_{{\bm{x}}_{k}}G}{\left\|G\right\|}-C\left(\frac{\nabla_{{\bm{x}}_{k}}G}{\left\|G\right\|^{3}}G\right)G^{\top}$
Using the sub-multiplicative property of the Frobenius norm and scaling property with high probability of $\left\|\nabla_{{\bm{x}}_{k}}G\right\|=\left\|\nabla_{{\bm{x}}_{k}}\nabla_{\theta}(y-f({\bm{x}}_{k}))^{2}\right\|_{F}^{2}=\tilde{O}(md)$ proved in Lemma D.6, we have that
$\left\|\nabla_{{\bm{x}}_{k}}{\bm{g}}_{\theta}\right\|_{F}^{2}\leq$ $2C^{2}\left\|\frac{\nabla_{{\bm{x}}_{k}}G}{\left\|G\right\|}\right\|_{F}^{2}+2C^{2}\left\|\frac{\nabla_{{\bm{x}}_{k}}G}{\left\|G\right\|^{3}}\right\|_{F}^{2}\left\|GG^{\top}\right\|_{F}^{2}$
$\leq$ $2C^{2}\frac{\tilde{O}(md)}{\left\|G\right\|^{2}}+2C^{2}\frac{\tilde{O}(md)}{\left\|G\right\|^{6}}\left\|GG^{T}\right\|_{F}^{2}$
$=$ $\frac{\tilde{O}(C^{2}md)}{\left\|G\right\|^{2}}(1+\left\|G\right\|^{-4}\left\|G\right\|^{4})$
$=$ $\frac{\tilde{O}(C^{2}md)}{\left\|G\right\|^{2}}.$
Therefore we have with high probability that
$\left\|\nabla_{{\bm{x}}}{\bm{g}}_{\theta}\right\|_{F}^{2}=$ $\sum_{i=1}^{B}\left\|\nabla_{{\bm{x}}_{i}}{\bm{g}}_{\theta}\right\|_{F}^{2}$
$=$ $\sum_{i=1}^{B}\frac{\tilde{O}(C^{2}md)}{\left\|G\right\|^{2}}$
$=$ $\frac{\tilde{O}(C^{2}mdB)}{\left\|G\right\|^{2}}$
Therefore our lower bound is given as
$R_{L}\geq\tilde{\Omega}(\sqrt{\frac{1}{B}\frac{B^{2}d^{2}}{\left\|\nabla_{{\bm{x}}}{\bm{g}}_{\theta}\right\|_{F}^{2}}\sigma^{2}})=\tilde{\Omega}(\sqrt{\frac{d}{m}}\frac{\left\|G\right\|}{C}),$
which is the desired result. ∎
E.3 Local Aggregation
In local aggregation, we calculate two updates of model parameters using two different batches of size $B$. In local aggregation, we use different learning rates for $a_{j}$ and ${\bm{w}}_{j}$ since their scalings are different. Specifically, we take $\eta_{a}=O(\frac{1}{m^{2}d})$ and $\eta_{w}=O(\frac{1}{d})$. Then $a_{j}^{(1)}=a_{j}-\eta_{a}\sum_{i=1}^{B}\nabla_{a_{j}}\ell$ and ${\bm{w}}_{j}^{(1)}={\bm{w}}_{j}-\eta_{w}\sum_{i=1}^{B}\nabla_{{\bm{w}}_{j}}\ell$ for all $j$. Denote $\theta^{(1)}=(a^{(1)},{\bm{w}}^{(1)})$ and we can similarly define $a^{(2)}$, ${\bm{w}}^{(2)}$ and $\theta^{(2)}$. Denote $\ell({\bm{x}}_{i},\theta,y_{i})=(y_{i}-f_{\theta}({\bm{x}}_{i}))$, the direct observation from the local aggregation gradients is (with $y_{j}$ omitted)
$a_{j}^{(2)}-a_{j}=-\eta_{a}\left(\sum_{i=1}^{B}\nabla_{a_{j}}\ell({\bm{x}}_{i},\theta)+\sum_{i=B+1}^{2B}\nabla_{a_{j}}\ell({\bm{x}}_{i},\theta)\right)$
and
${\bm{w}}_{j}^{(2)}-{\bm{w}}_{j}=-\eta_{w}\left(\sum_{i=1}^{B}\nabla_{{\bm{w}}_{j}}\ell({\bm{x}}_{i},\theta)+\sum_{i=B+1}^{2B}\nabla_{{\bm{w}}_{j}}\ell({\bm{x}}_{i},\theta)\right).$
In the setting of reconstruction attack, both of the learning rates are known so we can directly observe the gradients for two steps with a Gaussian noise.
$O(S)=\sum_{i=1}^{B}\nabla_{\theta}\ell({\bm{x}}_{i},\theta)+\sum_{i=B+1}^{2B}\nabla_{\theta^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right)+\epsilon,$ (30)
where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{md+d})$. We want to reconstruct all $2B$ data points from the update. Intuitively, this is approximately equivalent to a regular update with batch size $2B$ and learning rate $2\eta$, so the lower bound should remain unchanged. We now prove the bound rigorously.
Proposition E.2 (Full version of Proposition 3.2).
If satisfy assumption D.1, $m>B$ and the learning rate for the two layers $\eta_{w}=O(1)$ and $\eta_{a}=O(\frac{1}{m^{3/2}})$, then with high probability the lower bound of minimax risk $R_{L}$ with the observation Eq. (30) is of the scale $R_{L}\geq\tilde{\Omega}(\sigma\sqrt{\frac{d}{m}})$.
Proof.
Define
$g({\bm{x}}_{1:2B})=\sum_{i=1}^{B}\nabla_{\theta}\ell({\bm{x}}_{i},\theta)+\sum_{i=B+1}^{2B}\nabla_{\theta^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right),$
our lower bound is given as
$\frac{1}{2B}\frac{B^{2}d^{2}}{\sum_{i=1}^{2B}\left\|\nabla_{{\bm{x}}_{i}}g({\bm{x}}_{1:2B})\right\|^{2}}\sigma^{2}.$
Notice that all entries of $\nabla_{a}\ell({\bm{x}}_{i},\theta)$ are $\tilde{O}(1)$ and $\nabla_{w}\ell({\bm{x}}_{i},\theta)$ are $\tilde{O}(\frac{1}{m})$:
$\begin{cases}\nabla_{a_{j}}\ell=2\sum_{i=1}^{B}(y-f({\bm{x}}_{i}))\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})\\ \nabla_{{\bm{w}}_{j}}\ell=2\sum_{i=1}^{B}(y-f({\bm{x}}_{i}))a_{j}\sigma^{\prime}({\bm{w}}_{j}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}.\end{cases}$
Therefore taking $\eta_{a}\leq O(\frac{1}{B})$ and $\eta_{w}\leq O(\frac{m}{B})$ would make the updated $a_{j}^{(1)}=a_{j}-\eta_{a}\nabla_{a_{j}}\ell$ and ${\bm{w}}_{j}^{(1)}={\bm{w}}_{j}-\eta_{w}\nabla_{{\bm{w}}_{j}}\ell$ still satisfy scaling in Lemma D.5.
Therefore for $i\geq n+1$, with high probability
$\left\|\nabla_{{\bm{x}}_{i}}g({\bm{x}}_{1:2B})\right\|_{F}^{2}=\left\|\nabla_{{\bm{x}}_{i}}\nabla_{\theta^{(1)}}\ell\left({\bm{x}}_{i},\theta_{1}\right)\right\|_{F}^{2}=\tilde{O}(md).$
For $i\leq n$, with high probability
$\left\|\nabla_{{\bm{x}}_{i}}\nabla_{a_{j}}\ell\left({\bm{x}}_{i},\theta\right)\right\|^{2}=\tilde{O}(d),$ (31)
$\left\|\nabla_{{\bm{x}}_{i}}\nabla_{{\bm{w}}_{j}}\ell\left({\bm{x}}_{i},\theta\right)\right\|_{F}^{2}=\tilde{O}(\frac{d}{m^{2}}),$ (32)
$\left|\nabla_{a_{k}}\nabla_{a_{j}}\ell\left({\bm{x}}_{i},\theta\right)\right|=\tilde{O}(1)$ (33)
for any $j,k$,
$\left\|\nabla_{a_{j}}\nabla_{{\bm{w}}_{j}}\ell\left({\bm{x}}_{i},\theta\right)\right\|^{2}\leq\tilde{O}(1),$ (34)
$\left\|\nabla_{a_{k}}\nabla_{{\bm{w}}_{j}}\ell\left({\bm{x}}_{i},\theta\right)\right\|^{2}\leq\tilde{O}(\frac{1}{m^{2}})$ (35)
for $j\neq k$,
$\left\|\nabla_{{\bm{w}}_{j}}\nabla_{{\bm{w}}_{j}}\ell\left({\bm{x}}_{i},\theta\right)\right\|_{F}^{2}\leq\tilde{O}(\frac{1}{m^{2}})$ (36)
and
$\left\|\nabla_{{\bm{w}}_{k}}\nabla_{{\bm{w}}_{j}}\ell\left({\bm{x}}_{i},\theta\right)\right\|_{F}^{2}\leq\tilde{O}(\frac{1}{m^{4}})$ (37)
for $j\neq k$. Therefore, by Eq. (31) to (37) we have
$\left\|\nabla_{{\bm{x}}_{i}}g({\bm{x}}_{1:2B})\right\|^{2}=$ $\left\|\nabla_{{\bm{x}}_{i}}\nabla_{\theta^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)-\sum_{i=B+1}^{2B}\left(\eta_{w}\nabla_{{\bm{x}}_{i}}\nabla_{{\bm{w}}^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)\nabla_{{\bm{w}}^{(1)}}\nabla_{a^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right)\right.\right.$
$\left.+\eta_{w}\nabla_{{\bm{x}}_{i}}\nabla_{{\bm{w}}^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)\nabla_{{\bm{w}}^{(1)}}\nabla_{{\bm{w}}^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right)\right.$
$\left.+\eta_{a}\nabla_{{\bm{x}}_{i}}\nabla_{a^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)\nabla_{a^{(1)}}\nabla_{{\bm{w}}^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right)\right.$
$\left.\left.+\eta_{a}\nabla_{{\bm{x}}_{i}}\nabla_{a^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)\nabla_{a^{(1)}}\nabla_{a^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right)\right)\right\|_{F}^{2}$
$\lesssim$ $\left\|\nabla_{{\bm{x}}_{i}}\nabla_{\theta^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)\right\|_{F}^{2}+\sum_{i=1}^{B}\left(\eta_{w}\left\|\nabla_{{\bm{x}}_{i}}\nabla_{{\bm{w}}^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)\nabla_{{\bm{w}}^{(1)}}\nabla_{a^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right)\right\|_{F}^{2}\right.$
$+\eta_{w}\left\|\nabla_{{\bm{x}}_{i}}\nabla_{{\bm{w}}^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)\nabla_{{\bm{w}}^{(1)}}\nabla_{{\bm{w}}^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right)\right\|_{F}^{2}$
$+\eta_{a}\left\|\nabla_{{\bm{x}}_{i}}\nabla_{a^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)\nabla_{a^{(1)}}\nabla_{{\bm{w}}^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right)\right\|_{F}^{2}$
$+\left.\eta_{a}\left\|\nabla_{{\bm{x}}_{i}}\nabla_{a^{(1)}}\ell\left({\bm{x}}_{i},\theta\right)\nabla_{a^{(1)}}\nabla_{a^{(1)}}\ell\left({\bm{x}}_{i},\theta^{(1)}\right)\right\|_{F}^{2}\right)$
$\lesssim$ $\tilde{O}(md)+\sum_{i=1}^{B}\left(\eta_{w}^{2}m\tilde{O}(\frac{d}{m^{2}})(m\tilde{O}(1)+m(m-1)\tilde{O}(\frac{1}{m^{2}})\right.$
$+\eta_{w}^{2}m\tilde{O}(\frac{d}{m^{2}})(m\tilde{O}(\frac{1}{m^{2}})+m(m-1)\tilde{O}(\frac{1}{m^{4}}))$
$+\eta_{a}^{2}m\tilde{O}(d)(m\tilde{O}(1)+m(m-1)\tilde{O}(\frac{1}{m^{2}}))$
$\left.+\eta_{a}^{2}m\tilde{O}(d)m^{2}\tilde{O}(1)\right)$
$\leq$ $\tilde{O}(md)+\eta_{w}^{2}\tilde{O}(Bd)+\eta_{a}^{2}\tilde{O}(Bm^{3}d)$
$=$ $\tilde{O}(md)$
if $\eta_{w}=O(\sqrt{m}{B})$ and $\eta_{a}=O(\frac{1}{m\sqrt{B}})$.
Therefore the lower bound is
$R_{L}^{2}\geq\frac{1}{B}\frac{d^{2}B^{2}}{\sum_{i=1}^{2B}\left\|\nabla_{x_{i}}g(x_{1:2B})\right\|^{2}}\frac{1}{m^{2}}\sigma^{2}=\tilde{\Omega}(\frac{d}{m})\sigma^{2}.$
Thus, we have $R_{L}\geq\tilde{\Omega}(\sigma\sqrt{\frac{d}{m}})$. ∎
E.4 Gradient Pruning
Gradient pruning refers to changing small entries of the model gradient to zero before updating the model. If a model gradient was pruned, the entries corresponding to this term in the Jacobian matrix would become 0. Denote the Jacobian of pruned parameters’ gradient on the input ${\bm{x}}$ the matrix $J_{0}({\bm{x}})$ (which is part of the full Jacobian $\mathrm{tr}(J({\bm{x}})$), we will demonstrate in this section that gradient pruning has a lower bound at least as good as the regular bound in Proposition D.7.
Lemma E.3.
Under Assumption D.1, and for any constant $\gamma$, with high probability the lower bound of minimax risk $R_{L}$ with the observation
$P\left(\sum_{i=1}^{B}\nabla_{\theta}(y-f({\bm{x}}_{i}))^{2}\right)+{\bm{\epsilon}}$
is of scale $R_{L}\geq\tilde{\Omega}(\sigma\sqrt{\frac{d}{m(1-\hat{p})}})$, where $P(\mathbf{a})$ is the function that for all $i=1,\dots,\dim(\mathbf{a})$,
$P_{i}(\mathbf{a})=\begin{cases}a_{i}&\text{if }|a_{i}|\geq\gamma\\ 0&\text{if }|a_{i}|<\gamma\end{cases},$
and
$\hat{p}=\frac{\left\|J_{0}({\bm{x}})\right\|_{F}^{2}}{\left\|J({\bm{x}})\right\|_{F}^{2}}.$
Proof.
As the points that are not differentiable after applying this defense method are of zero measure, our analysis could ignore these cases.
The Jacobian $\nabla_{\bm{x}}P\left(\sum_{i=1}^{B}\nabla_{\theta}(y-f({\bm{x}}_{i}))^{2}\right)$ is $\nabla_{\bm{x}}\sum_{i=1}^{B}\nabla_{\theta}(y-f({\bm{x}}_{i}))^{2}$ with the columns corresponding to columns in $J_{0}$ being set to 0. Our lower bound is given as
$R_{L}\geq\tilde{\Omega}(\sigma\sqrt{\frac{1}{B}\frac{d^{2}B^{2}}{\left\|J({\bm{x}})\right\|_{F}^{2}-\left\|J_{0}{{\bm{x}}}\right\|_{F}^{2}}})=\tilde{\Omega}(\sigma\sqrt{\frac{1}{B}\frac{d^{2}B^{2}}{\left\|J({\bm{x}})\right\|_{F}^{2}}\frac{1}{1-\hat{p}}}).$
Which gives us the desired result. ∎
Remark E.4.
Since entries of $G_{{\bm{w}}_{j}}$ scale smaller compared to $G_{a_{j}}$, gradients from the first layer easier to be pruned. Therefore until the ratio of pruned gradients reaches $\frac{d}{d+1}$, the pruned gradients gives a small $\hat{p}$ and does not change the lower bound much. But higher than this ratio we start to prune gradients of $a_{j}$, so we predict that the bound will greatly increase after the ratio reaches $\frac{d}{d+1}$.
Appendix F Additional experimental details for our proposed attack
@@
The architecture of our deep image prior network is structured on the foundation of PixelCNN++ as per [Salimans et al., 2017], and it utilizes a U-Net as depicted by [Ronneberger et al., 2015], which is built upon a Wide-ResNet according to [Zagoruyko and Komodakis, 2016], To simplify the implementation, we substituted weight normalization [Salimans and Kingma, 2016] with group normalization [Wu and He, 2018]. The models operating on $32\times 32$ utilize four distinct feature map resolutions, ranging from $32\times 32$ to $4\times 4$. Each resolution level of the U-Net comprises two convolutional residual blocks along with self-attention blocks situated at the $16\times 16$ resolution between the convolutional blocks. When available, the reconstructed features are incorporated into each residual block.
@@
We adopt ResNet-18 [He et al., 2016] architecture as our backbone network. The network is pre-trained for one epoch for better results of gradient inversion. Note that this architecture results in a more challenging setting given its minimal information retention capacity and is observed to cause a slight drop in quantitative performance and large variance [Jeon et al., 2021]. We further add an extra linear layer before the classification head for feature reconstruction. In order to satisfy the concentration bound $\tilde{O}(B\sqrt{\frac{d}{m}})$ of feature reconstruction, this linear layer has $512\times 32=16384$ neurons.
@@
The attack minimize the objective function given in Eq. 4. We search $\alpha_{\rm TV}$ in $\{0,0.001,0.005,0.01,0.05,0.1,0.5\}$ and $\alpha_{\rm BN}$ in $\{0,0.0005,0.001,0.01,0.05,0.01\}$ for attack without defense, and apply the best choices $\alpha_{\rm TV}=0.01$ and $\alpha_{\rm BN}=0.001$ for all defenses (we did not tune $\alpha_{\rm TV}$ and $\alpha_{\rm BN}$ for each defense as we observe the proposed attack method is robust in terms of these two hyper-parameters). We optimize the attack for 10,000 iterations using Adam [Kingma and Ba, 2014], with the initial learning rate set to $0.00001$. We decay the learning rate by a factor of $0.1$ at $3/8$, $5/8$, and $7/8$ of the optimization. For feature matching loss strength $\alpha_{f}$, we search in $\{0.01,0.05,0.1,0.5\}$ for each defense and find that $0.1$ is the optimal choice.
@@
Zhu et al., 2019, Geiping et al., 2020 suggests that small batch size is vital for the success of the attack. We therefore intentionally evaluate the attack with three small batch sizes $2$, $4$, and $8$ to test the upper bound of privacy leakage, and the minimum (and unrealistic) batch size $1$, and two small but realistic batch sizes $16$ and $32$.
@@
The feature reconstruction method suggested by Wang et al., 2023 always produces features that are normalized and may have a sign opposite to the actual features. To compare the similarity between the original and the reconstructed features, we use cosine similarity, as it doesn’t change with feature magnitude. We square this similarity to avoid any issues with mismatched signs.
Appendix G Additional experimental details for other methods
For GradientInversion [Geiping et al., 2020], we adopted the code from [Huang et al., 2021], and the hyperparameters providedhttps://github.com/Princeton-SysML/GradAttack/. For Rob The Fed [Fowl et al., 2021], we follow strictly the setting mentioned in their paper and the code§§§https://github.com/lhfowl/robbing_the_fed, and introduced defenses to the gradients. For CPA [Kariyappa et al., 2023], since the settings and model architectures used in their paper are very different from ours, we incorporate their algorithm for feature recovery into our settings. i.e. using the latent features recovered by CPA rather than our proposed method. We perform hyperparameter selection through grid searches, but due to limited time, the hyperparameter searching process may not be comprehensive. But this should not significantly deviate our main goal – comparison of the effectiveness of different defense methods, (not the attack algorithms).
Appendix H Hardware and software
We run most of the experiments on clusters using NVIDIA V100s. All experiments are implemented using PyTorch library. Overall we estimated that a total of 800 GPU hours were consumed.
Appendix I Additional experimental results